OpenStudy (anonymous):

Algebra 1 help please medal will be awarded!

OpenStudy (anonymous):

I do not know 1 please help

jimthompson5910 (jim_thompson5910):

Surface Area: Use the formula SA = 2*(LW + LH + WH)

jimthompson5910 (jim_thompson5910):

Volume: Use the formula V = L*W*H

OpenStudy (anonymous):

okay so #2 is 64,288 ft^3 right

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

okay good. i am still confused about 1 though

jimthompson5910 (jim_thompson5910):

SA = 2*(LW + LH + WH) SA =2*(41*32+32*49+41*49) SA = ??

OpenStudy (anonymous):

Okay so #1 is 7,154 ft^2 right?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

try it again

OpenStudy (anonymous):

okay hold on

jimthompson5910 (jim_thompson5910):

use a calculator

jimthompson5910 (jim_thompson5910):

if you don't have one, use google as a calculator

OpenStudy (anonymous):

oh....thank you

jimthompson5910 (jim_thompson5910):

np